# Classification of fractional, singular Yamabe metrics on a twice-punctured sphere I

### v1

@@ 3.2. Sharp supremum estimate
This is an adaptation of the proof of [Azahara_Sergio_Ruiz, Proposition 2.5]. Following their notation, for any $L>0$ we define
@@
and let $\phi_{1}\in H^{s}(\mathbf{R})\backslash\{0\}$ be the associated eigenfunction. They show in [Azahara_Sergio_Ruiz, Lemma 2.4] that the eigenvalue is positive for each $L>0$ and that $\lim_{L\rightarrow\infty}\lambda_{1}(L)=0$, as well as the fact that the eigenfunction is a positive solution to
@@
$\left(\frac{n+2s}{n-2s}\right)\hat{c}_{n,s}^{\frac{4s}{n-2s}}(\cosh L_{0})^{-2s}=\frac{1}{2}$
@@
so we can conclude that $P_{s}(w)\geq 0$ and use the maximum principle (see [Azahara_Sergio_Ruiz, Proposition 2.2]) as soon as we know $\lambda_{1}(L_{0})\leq\frac{c_{n,s}}{2}$. However, this inequality follows from [Azahara_Sergio_Ruiz, Eq. (2.4)] and the subsequent scaling law. ∎
@@
In some sense, this estimate is similar to the height estimate of a constant mean curvature graph in [KKS, Lemma 1.7].
@@
Then, by the classification result in [MR2200258, Theorem 1.1], we know that any solution $v\in\mathcal{C}^{2}(\mathbf{R})$ with fast-decay $\lim_{|t|\to\infty}v(t)=0$ must be a translation of the spherical solution $v_{\circ}$. Since $v\not\equiv v_{\circ}$, there holds
@@ 4.1. Local classification
Proof of Theorem 1 (local part).
Assume there exist a sequence $\{s_{k}\}_{k\in\mathbf{N}}\subset(0,1)$ of fractional powers such that $s_{k}\nearrow 1$ as $k\to\infty$ and a sequence of positive smooth solutions $\{v_{k}\}_{k\in\mathbf{N}}\subset\mathcal{C}^{2}(\mathbf{R})$ to ($\widehat{\mathcal{Q}}_{n,s}$), that is
@@
$\|v_{k}\|_{L^{\infty}(\mathbf{R})}=\|\overline{v}_{k}\|_{L^{\infty}(\mathbf{R})}\quad{\rm and}\quad v_{k}(0)=\|v_{k}\|_{L^{\infty}(\mathbf{R})}=\|\overline{v}_{k}\|_{L^{\infty}(\mathbf{R})}=\overline{v}_{k}(0).$
@@
Fix $L>0$ and define a normalized sequence $\{\psi_{k}\}_{k\in\mathbb{N}}\subset\mathcal{C}^{2}(\mathbf{R})$ such that
@@
Claim 1: For each fixed $L>0$, there exists $\psi_{\infty}\in\mathcal{C}^{2,\alpha}([-L,L])$ such that
$\lim_{k\to\infty}\|\psi_{k}-\psi_{\infty}\|_{\mathcal{C}^{0}([-L,L])}=0.$
Proof of Claim 1: We argue by contradiction and, without loss of generality, work along a subsequence that preserves convergence of all relevant quantities. Recall that each $\overline{v}_{k}$ is a Delaunay-type solution to ($\widehat{\mathcal{Q}}_{n,s}$) corresponding to the same necksize normalization
@@
$\mathcal{L}_{*}(\psi):=-\psi^{\prime\prime}+c_{n,1}\,\psi-\tfrac{n+2}{n-2}\,c_{n,1}\,\overline{v}_{*}^{\frac{4}{n-2}}\psi\quad{\rm for}\quad\psi\in\mathcal{C}^{2}(\mathbf{R}).$ (23)
This operator is precisely the local ($s=1$) limit of the family $\mathcal{L}_{k}$, since $P_{s_{k}}\to-\partial_{t}^{2}$ in the strong resolvent sense as $s_{k}\to 1^{-}$ and $\overline{v}_{k}\to\overline{v}_{*}$ in $\mathcal{C}^{2}_{\mathrm{loc}}$.
@@
From this, we can apply the Arzelà–Ascoli theorem to find function $\psi_{\infty}\in\mathcal{C}^{2,\alpha}([-L,L])$ such that, up to subsequence, $\psi_{k}\rightarrow\psi_{\infty}$ in $\mathcal{C}^{2,\alpha}([-L,L])$ satisfying
$\begin{cases}&-\psi_{\infty}^{{}^{\prime\prime}}+\left(\frac{n-2}{2}\right)^{2}\psi_{\infty}-V_{\infty}(\zeta)\psi_{\infty}=0\quad\text{in}\quad(-L,L)\\ &\psi_{\infty}(0)=1.\end{cases}$ (25)
@@
In view of (23) and the initial conditions $\psi_{\infty}(0)=\psi_{\infty}^{\prime}(0)=0$, the uniqueness theorem for ODEs implies $\psi_{\infty}\equiv 0$ on [L,L][-L,L]. However, by normalization we have
$\|\psi_{\infty}\|_{L^{\infty}([-L,L])}=\lim_{k\to\infty}\|\psi_{k}\|_{L^{\infty}([-L,L])}=1,$
which is a contradiction. This completes the proof of the claim.
Indeed, we suppose that the claim is not valid and then derive a contradiction. Initially, by construction, one has $\|\psi_{k}\|_{L^{\infty}([-L,L])}=1$ for each $k\in\mathbf{N}$. This allows us to apply Arzelà–Ascoli to find $\psi_{\infty}\in\mathcal{C}^{2,\alpha}([-L,L])$ for some $\alpha\in(0,1)$ such that $\|\psi_{k}-\psi_{\infty}\|_{\mathcal{C}^{2,\alpha}([-L,L])}\to 0$ as $k\to\infty$ up to subsequence, and satisfies
$\mathcal{L}_{*}(\psi_{\infty})=0,$
where $\mathcal{L}_{*}$ is the linearized operator about $\overline{v}_{*}$. However, this last equation states that $\psi_{\infty}$ satisfies a homogeneous, second-order ODE, which, combined with the initial conditions $\psi_{\infty}(0)=0=\psi_{\infty}^{\prime}(0)$ implies $\psi_{\infty}\equiv 0$. This contradicts the normalization
$\|\psi_{\infty}\|_{L^{\infty}([-L,L])}=\lim_{k\rightarrow\infty}\|\psi_{k}\|_{L^{\infty}([-L,L])}=1,$
and concludes the proof of the claim. ∎
@@ 4.2. Global classification
We now pass from the local convergence to global convergence. The key inputs are a weighted norm that matches the exponential and an ODE barrier lemma for noncompact recenterings.
Before proving global classification, we record the ODE barrier that excludes fast left decay for global limit profiles.
@@
Let us define $w(y):=e^{-\sigma_{*}\zeta}\psi(\zeta)$ for $\zeta\leq 0$. A direct computation gives us
@@
Integrating once more, and using $w(\zeta)\to 0$ as $\zeta\to$-$\infty$, we get the Volterra-type identity
$w(\zeta)=-\frac{1}{2\sigma_{*}}\int_{-\infty}^{y}\big(1-e^{-2\sigma_{*}(\zeta-\tau)}\big)V(\tau)w(\tau)d\tau.$
@@
$P_{s}(\phi)(t)-c_{n,s}\phi(t)+\ V_{s}(t)\phi(t)=\varphi_{s}(t)\quad{\rm in}\quad I$ (28)
@@
This is a direct consequence of the regularity results by Silvestre [MR2244602, Theorem 5.1] and Ros-Oton and Serra [MR3482695, Theorem 1.1] under the verification that the log-cylindrical operator $\mathbb{P}_{s}=P_{s}+c_{n,s}$ on the right-hand side of (28) satisfies the ellipticity conditions [MR2244602, Eq. (2.2), pg. 1158] and [MR3482695, Eq. (1.2), pg. 8676]. Here, we notice that by [MR3542618, Theorem 1.1] that there exists $s_{0}\in(\frac{1}{2},1)$ such that the regularity constant $C>0$ does not depend on $s\in(s_{0},1]$. ∎
@@
with $0<c_{1}<V_{s_{k}}(t)<c_{2}$ for $s_{k}\in(0,1)$, where to get the initial conditions (21) we used (21).
@@
Setting $\rho=t+\tau$ and using [MR3694655, Lemmas 2.5 and 2.6], one knows that there exist $\varepsilon,M\in\mathbf{R}$ with $0\ll\varepsilon\ll 1$ small and $M\gg 1$ large such that the kernel of $P_{s}$ acting on radial functions satisfies
$K_{s}(\rho)\sim\begin{cases}\varsigma_{n,s}|\rho|^{-(1-2s)},\quad{\rm if}\quad&0<|\rho|<\varepsilon\\ \varsigma_{n,s},\quad{\rm if}\quad&0\leq{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\varepsilon<}|\rho|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\leq M}\leq\infty\\ \varsigma_{n,s}e^{-\frac{n+2s}{s}|\rho|,}\quad{\rm if}\quad&|\rho|>M,\end{cases}$
@@
It remains to estimate the integral terms $I_{1}$, $I_{2}$, and $I_{3}$. First, by using the regularity in Lemma 11 and the parity of the kernel, we can Taylor expand $\phi_{k}(t-\rho)$ to get
$I_{1}$ $=\int_{-\varepsilon}^{\varepsilon}|\rho|^{-(1+2s_{k})}(\phi_{k}(t-\rho)-\phi_{k}(t))\,d\rho$ (34)
$=\int_{-\varepsilon}^{\varepsilon}|\rho|^{-(1+2s)}\left[-\phi_{k}^{\prime}(t)(-\rho)+\phi_{k}^{\prime\prime}(t)\frac{\rho^{2}}{2}+\mathcal{O}(\rho^{3})\right]\,d\rho=\frac{\varepsilon^{2-2s_{k}}}{2-2s_{k}}\phi_{k}^{\prime\prime}(t)+\mathcal{O}(\varepsilon^{2})\leq\frac{C_{1}}{2-2s_{k}}\phi_{k}^{\prime\prime}(t)$
for some $C_{1}=C_{1}(n,\varepsilon)>0$. Second, using the bound (32), it is straightforward to see that
@@
$P_{s_{k}}(\phi_{k})\sim(1-s_{k})\left[\frac{C_{1}}{2(1-s_{k})}\phi_{k}^{\prime\prime}+C_{2}+C_{3}\right]\sim\frac{1}{2}\phi_{k}^{\prime\prime}\sim\phi_{k}^{\prime\prime}\quad\text{as}\quad s_{k}\to{1^{-}}.$
@@
$\begin{cases}-\phi_{\infty}^{{}^{\prime\prime}}+c_{n,1}\phi_{\infty}-V_{\infty}(t)\phi_{\infty}=0\quad\text{in}\quad\mathbf{R}\\ \phi_{\infty}(0)=\phi_{\infty}^{{}^{\prime}}(0).\end{cases}$ (37)
@@
and there exists a constant $c_{2}>0$ such that $0\leq V_{\infty}(t)\leq c_{2}$. Finally, using (38) we can apply the Lemma 10 to (39) and show that the decaying rate as $\zeta\to$-$\infty$ is too large. Thus, $\psi_{\infty}\equiv 0$, which is a contradiction to $\psi_{\infty}(0)=1$, finishing the proof. ∎
@@ 6.1. Extending the range of the fractional parameter
It is quite natural to ask whether our classification result remains valid when the order of the conformal operator exceeds one. In this subsection we discuss our work in progress [our_followup] to classify conformally flat, constant $Q_{s}$-curvature metrics on a twice-punctured sphere in the case that $s\in(1,1+\delta)$, where $\delta$ is sufficiently small.
The first complication in this classification lies in the description of the Delaunay solutions of order $s$ when $s>1$. Indeed, the integral operator $P_{s}$ as given in (9) is not well-defined in the case $s>1$, so one needs an alternative formulation of the problem. Jin and Xiong [arxiv:1901.01678] give a dual formulation of the problem, starting with the equation
@@
The larger complication is that the passage from local convergence to global convergence, as described in our Section 4.2 above, relies on a sharp estimate of the growth rate of solutions to the linearized equation. In the regime $s>1$ these estimates are not currently available, and in particular we do not have an analog of Lemma 10 in this setting. Deriving such estimates associated to the linearization of the dual operator will occupy a large portion of [our_followup].
@@ 6.2. Informed speculation
As a direct consequence of our classification, one can rapidly prove several related results. First of all, one can combine the nondegenerate with the slide-back technique in the paper by Korevaar, Mazzeo, Pacard and Schoen [MR1666838] to obtain refined asymptotics of solutions with isolated singularities. Second, we expect that one can adapt the proof of compactness in the moduli space setting, as in the paper by Pollack [MR1266101]. We only sketch the ideas of both of these proofs and leave the details for future projects.

### v2

@@ 3.2. Sharp supremum estimate
This is an adaptation of the proof of [14, Proposition 2.5]. Following their notation, for any $L>0$ we define
@@
and let $\phi_{1}\in H^{s}(\mathbf{R})\backslash\{0\}$ be the associated eigenfunction. They show in [14, Lemma 2.4] that the eigenvalue is positive for each $L>0$ and that $\lim_{L\rightarrow\infty}\lambda_{1}(L)=0$, as well as the fact that the eigenfunction is a positive solution to
@@
$\left(\frac{n+2s}{n-2s}\right)\hat{c}_{n,s}^{\frac{4s}{n-2s}}(\cosh L_{0})^{-2s}=\frac{3}{2}$
@@
so we can conclude that $P_{s}(w)\geq 0$ and use the maximum principle (see [14, Proposition 2.2]) as soon as we know $\lambda_{1}(L_{0})\leq\frac{c_{n,s}}{2}$. However, this inequality follows from [14, Eq. (2.4)] and the subsequent scaling law. ∎
@@
In some sense, this estimate is similar to the height estimate of a constant mean curvature graph in [25, Lemma 1.7].
@@
Then, by the classification result in [12, Theorem 1.1], we know that any solution $v\in\mathcal{C}^{2}(\mathbf{R})$ with fast-decay $\lim_{|t|\to\infty}v(t)=0$ must be a translation of the spherical solution $v_{\circ}$. Since $v\not\equiv v_{\circ}$, there holds
@@ 4.1. Local classification
Proposition 9.
Let $n\in\mathbf{N}$ and $s\in(0,1)$ with $n\geq 3$. For each fixed $L>0$, there exists $\delta>0$ such that if $1-\delta<s<1$ and $v:\mathbf{R}\rightarrow(0,\infty)$ satisfies ($\widehat{\mathcal{Q}}_{n,s}$), then the restriction of $v$ to [L,L][-L,L] is a Delaunay solution.
Proof.
We prove this proposition by contradiction. Assume there exist a sequence $\{s_{k}\}_{k\in\mathbf{N}}\subset(0,1)$ of fractional powers such that $s_{k}\nearrow 1$ as $k\to\infty$ and a sequence of positive smooth solutions $\{v_{k}\}_{k\in\mathbf{N}}\subset\mathcal{C}^{2}(\mathbf{R})$ to ($\widehat{\mathcal{Q}}_{n,s}$), that is
@@
$\|v_{k}\|_{L^{\infty}([-L,L])}=\|\overline{v}_{k}\|_{L^{\infty}([-L,L])}\quad{\rm and}\quad v_{k}(0)=\|v_{k}\|_{L^{\infty}([-L,L])}=\|\overline{v}_{k}\|_{L^{\infty}([-L,L])}=\overline{v}_{k}(0).$
@@
Define a normalized sequence $\{\psi_{k}\}_{k\in\mathbb{N}}\subset\mathcal{C}^{2}(\mathbf{R})$ such that
@@
Recall that each $\overline{v}_{k}$ is a Delaunay-type solution to ($\widehat{\mathcal{Q}}_{n,s}$) corresponding to the same necksize normalization
@@
$\mathcal{L}_{*}(\psi):=-\psi^{\prime\prime}+c_{n,1}\,\psi-\left(\tfrac{n+2}{n-2}\right)\,c_{n,1}\,\overline{v}_{*}^{\frac{4}{n-2}}\psi\quad{\rm for}\quad\psi\in\mathcal{C}^{2}(\mathbf{R}).$ (24)
This operator is precisely the local ($s=1$) limit of the family $\mathcal{L}_{k}$, since $P_{s_{k}}\to-\partial_{t}^{2}$ in the strong resolvent sense as $s_{k}\to 1^{-}$ (see the estimates (4.2), (36) and (37) below) and $\overline{v}_{k}\to\overline{v}_{*}$ in $\mathcal{C}^{2}_{\mathrm{loc}}$.
@@
From this, we can apply the Arzelà–Ascoli theorem to find a function $\psi_{\infty}\in\mathcal{C}^{2,\alpha}([-L,L])$ such that, up to subsequence, $\psi_{k}\rightarrow\psi_{\infty}$ in $\mathcal{C}^{2,\alpha}([-L,L])$ satisfying
$\begin{cases}&-\psi_{\infty}^{{}^{\prime\prime}}+\left(\tfrac{n-2}{2}\right)^{2}\psi_{\infty}-\left(\tfrac{n(n+2)}{4}\right)v_{*}^{\frac{4}{n-2}}\psi_{\infty}=0\quad\text{in}\quad(-L,L)\\ &\psi_{\infty}(0)=0,\quad\psi_{\infty}^{\prime}(0)=0.\end{cases}$ (26)
@@
In view of (24) and the initial conditions $\psi_{\infty}(0)=\psi_{\infty}^{\prime}(0)=0$, the uniqueness theorem for ODEs implies $\psi_{\infty}\equiv 0$ on [L,L][-L,L], which in turn contradicts the normalization $\|\psi_{k}\|_{L^{\infty}([-L,L])}=1$ for each $k$. ∎
@@ 4.2. Global classification
In this section, we complete the proof of Theorem 1 by showing the convergence of $\{\psi_{k}\}$ to $\psi_{\infty}$ is global, in $L^{\infty}(\mathbf{R})$. In other words, we show that one can choose the same $\delta$ for each $L$ in Proposition 9.
Before proving global convergence, we record the ODE barrier that excludes fast left decay for global limit profiles.
@@
Let us define $w(\zeta):=e^{-\sigma_{*}\zeta}\psi(\zeta)$ for $\zeta\leq 0$. A direct computation gives us
@@
Integrating by parts and using the Fundamental Theorem of Calculus along with the boundary condition $w(\zeta)\to 0$ as $\zeta\to$-$\infty$, we get the Volterra-type identity
$w(\zeta)=-\frac{1}{2\sigma_{*}}\int_{-\infty}^{\zeta}\big(1-e^{-2\sigma_{*}(\zeta-\tau)}\big)V(\tau)w(\tau)d\tau.$
@@
$P_{s}(\phi)(t)+c_{n,s}\phi(t)-\ V_{s}(t)\phi(t)=\varphi_{s}(t)\quad{\rm in}\quad I$ (29)
@@
This is a direct consequence of the regularity results by Silvestre [34, Theorem 5.1] and Ros-Oton and Serra [31, Theorem 1.1] under the verification that the log-cylindrical operator $\mathbb{P}_{s}=P_{s}+c_{n,s}$ on the right-hand side of (29) satisfies the ellipticity conditions [34, Eq. (2.2), pg. 1158] and [31, Eq. (1.2), pg. 8676]. Here, we notice that by [22, Theorem 1.1] there exists $s_{0}\in(\frac{1}{2},1)$ such that the regularity constant $C>0$ does not depend on $s\in(s_{0},1]$. ∎
@@
with $0<c_{1}<V_{s_{k}}(t)<c_{2}$ for $s_{k}\in(0,1)$, where we used (22).
@@
Setting $\rho=t-\tau$ and using [16, Lemmas 2.5 and 2.6], one knows that there exist $\varepsilon,M\in\mathbf{R}$ with $0\ll\varepsilon\ll 1$ small and $M\gg 1$ large such that the kernel of $P_{s}$ acting on radial functions satisfies
$K_{s}(\rho)\sim\begin{cases}\varsigma_{n,s}|\rho|^{-(1+2s)},\quad{\rm if}\quad&0<|\rho|<\varepsilon\\ \varsigma_{n,s},\quad{\rm if}\quad&0\leq\varepsilon<|\rho|\leq M<\infty\\ \varsigma_{n,s}e^{-\frac{n+2s}{s}|\rho|,}\quad{\rm if}\quad&|\rho|>M,\end{cases}$
@@
It remains to estimate the integral terms $I_{1}$, $I_{2}$, and $I_{3}$. First, by using the regularity in Lemma 11 and the parity of the kernel, we can Taylor expand $\phi_{k}(\tau)$ to get
$I_{1}$ $=\int_{|t-\tau|<\varepsilon}(\phi_{k}(t)-\phi_{k}(\tau))|t-\tau|^{-(1+2s_{k})}\,d\tau$
$=\int_{|t-\tau|<\varepsilon}\left[-\phi^{\prime}_{k}(t)(\tau-t)-\phi^{\prime\prime}_{k}(t)\frac{(\tau-t)^{2}}{2}+\mathcal{O}(|\tau-t|^{3})\right]|t-\tau|^{-(1+2s_{k})}\,d\tau$
$=\int_{-\varepsilon}^{\varepsilon}-|\rho|^{-(1+2s)}\left[\phi_{k}^{\prime}(t)\rho+\phi_{k}^{\prime\prime}(t)\frac{\rho^{2}}{2}+\mathcal{O}(|\rho|^{3})\right]\,d\rho$
$=-\frac{\varepsilon^{2-2s_{k}}}{2-2s_{k}}\phi_{k}^{\prime\prime}(t)+\mathcal{O}(\varepsilon^{2})=-\frac{1}{2(1-s_{k})}\phi_{k}^{\prime\prime}(t)+\mathcal{O}(\varepsilon^{2})\quad{\rm as}\quad s_{k}\to 1^{-}$ (35)
Second, using the bound (33), it is straightforward to see that
@@
$P_{s_{k}}(\phi_{k})\sim(1-s_{k})\left[\frac{1}{2(1-s_{k})}\phi_{k}^{\prime\prime}+C_{2}+C_{3}\right]\sim-\frac{1}{2}\phi_{k}^{\prime\prime}\sim-\phi_{k}^{\prime\prime}\quad\text{as}\quad s_{k}\to{1^{-}}.$
@@
$\begin{cases}\phi_{\infty}^{{}^{\prime\prime}}+c_{n,1}\phi_{\infty}-V_{\infty}(t)\phi_{\infty}=0\quad\text{in}\quad\mathbf{R}\\ \phi_{\infty}(0)=\phi_{\infty}^{{}^{\prime}}(0)=0,\end{cases}$ (38)
where $V_{\infty}(t)=\frac{n(n+2)}{4}(v_{*}(t))^{\frac{4}{n-2}}$ and $v_{*}=\lim_{k\rightarrow\infty}\overline{v}_{k}$ is the limit Delaunay solution.
@@
and there exists a constant $c_{2}>0$ such that $0\leq V_{\infty}(t)\leq c_{2}$. Finally, using (39) we can apply the Lemma 10 and the following remark to (40) and show that the decaying rate as $\zeta\to$-$\infty$ is too large. Thus, $\psi_{\infty}\equiv 0$, which is a contradiction to $\psi_{\infty}(0)=1$, finishing the proof. ∎
@@ 6.1. Extending the range of the fractional parameter
It is quite natural to ask whether our classification result remains valid when the order of the conformal operator exceeds one. In this subsection we discuss our work in progress [1] to classify conformally flat, constant $Q_{s}$-curvature metrics on a twice-punctured sphere in the case that $s\in(1,1+\delta)$, where $\delta$ is sufficiently small.
The first complication in this classification lies in the description of the Delaunay solutions of order $s$ when $s>1$. Indeed, the integral operator $P_{s}$ as given in (9) is not well-defined in the case $s>1$, so one needs an alternative formulation of the problem. Jin and Xiong [23] give a dual formulation of the problem, starting with the equation
@@
The larger complication is that the passage from local convergence to global convergence, as described in our Section 4.2 above, relies on a sharp estimate of the growth rate of solutions to the linearized equation. In the regime $s>1$ these estimates are not currently available, and in particular we do not have an analog of Lemma 10 in this setting. Deriving such estimates associated to the linearization of the dual operator will occupy a large portion of [1].
@@ 6.2. Informed speculation
As a direct consequence of our classification, one can rapidly prove several related results. First of all, one can combine the nondegenerate with the slide-back technique in the paper by Korevaar, Mazzeo, Pacard and Schoen [24] to obtain refined asymptotics of solutions with isolated singularities. Second, we expect that one can adapt the proof of compactness in the moduli space setting, as in the paper by Pollack [29]. We only sketch the ideas of both of these proofs and leave the details for future projects.